peeka1k:

In order for the data in the table to represent a linear function with a rate of change of +5, what must be the value of a? a = 3 a = 8 a = 18 a = 33

jhonyy9:

so how you think then you see this ? a=3 a=3+5=8 a=8+2*5 = 18 a=18+3*5=33 a=?